OpenStudy (anonymous):

solve algebraically (x+2)^2+(y-16)^2=1 y=-4x+8

OpenStudy (anonymous):

subtitue y=-4x+8 in the 1st equation ===> (x+3)^2 + (-4x+8-16)^2=1 ===> x^2+6x+9+(-4x-8)^2=1 ==> x^2+6x+9-(4x+8)^2=1 ===> x^2+6x+9-(16x^2+64x+64) =1 ===> x^2+6x+9-16x^2-64x-64 =1 ===> -15x^2 -58x -56=0 ===> -(15x^2 +58x +56)=0 ===> using the quadratic formula ==> x=-2 ==> y=16, or x=-28/15 ==> y=8/15

OpenStudy (anonymous):

wish it would be right becasuse alot of computaions